Please help us figure out how to get the formula.
Marisol is painting on a piece of canvas that has an area of 180 square inches. The length of the painting is 1 1/4 times the width. What are the dimensions of the painting?

Answer by reviewermath(1029)

Marisol is painting on a piece of canvas that has an area of 180 square inches. The length of the painting is 1 1/4 times the width. What are the dimensions of the painting?
Solution:
The mixed number 1 1/4 is equal to the improper fraction 5/4.
Let x = width
so (5/4)x = length
The area of a rectangle is equal to the product of length and width.
[(5/4)(x)](x) = 180
(5/4)(x^2) = 180
Multiply both sides by 4/5
x^2 = 144
Get the positive square root of both sides
x = 12
so (5/4)(x) = 15
Therefore, the width is 12 inches and the length is 15 inches.